The sum of three consecutive odd numbers is 51

Answer by ikleyn(52781)

.

The sum of three consecutive odd numbers is 51

                    ~~~~~~~~~~~~~~~


Let the numbers be  (x-2),  x,  (x+2).


Then the sum is   (x-2) + x + (x+2) = 3x.


The equation is           3x = 51.


The solution is     x = 51/3 = 17.


The  ANSWER  is:   the numbers are  15,  17,  19.



The lesson to learn:               the sum of three consecutive odd integer numbers is thrice the middle number.


The other lesson to learn:     if the sum of three consecutive odd integer numbers is given,  then the middle number is one third of the sum.



                    Happy learning   (!)